OpenStudy (ttp):

Which of the following ordered pairs is a solution to the system of linear equations below? Three x plus y equals nine. Two x minus two y equals fourteen. (2, 3) (3, 2) Four, negative three. Negative three, four.

OpenStudy (esshotwired):

3x+y=9 2x-2y=14 write it out so it is easier to visualize

OpenStudy (the_fizicx99):

Didn't you solve a 3 variable system of equations? Anyways, 3x + y = 9 2x - 2y = 14 3x *(-2) + y*(-2) = 9*(-2) 2x*(3) -y*(3) = 14*3 -6x -2y = -18 6x -6y = 42 0 -8y = 24 /-8 /-8 y = -3 Now you solve for X. Plug in Y.

OpenStudy (ttp):

3x+y=9 2x-2y=14

OpenStudy (the_fizicx99):

Can you do it? If you want I can guide you along it. :)

OpenStudy (ttp):

I'm confused :)

OpenStudy (the_fizicx99):

Sorry, k. So when you plug in Y for your first equation you should get. 3x + (-3) = 9 A positive and a negative = negative So this is: 3x -3 = 9 add 3 to both sides to leave x alone on the left.

OpenStudy (ttp):

x=4 right?

OpenStudy (the_fizicx99):

Yup!! So now plug in x and y and it should give you 9 = 9 When it does those are your coordinates!!

OpenStudy (ttp):

Where do i plug in what?

OpenStudy (the_fizicx99):

xD you keep changing your picture lol um plug in x and y 3(4) + (-3) = 9

OpenStudy (ttp):

*proves

OpenStudy (the_fizicx99):

Fake?

OpenStudy (ttp):

I said It proves I'm not fake

OpenStudy (ttp):

Cause alot of people say im too pretty to be real lol

OpenStudy (the_fizicx99):

That means 2 things lol xDDD and yea, um 3*4 - 3 = 9 3*4 = 12 12 - 3 = 3 12-3 = 9 9 = 9 Solution!! It works. You plug it in for the other equation and see if its true

OpenStudy (the_fizicx99):

12 - 3 = 9 *

OpenStudy (ttp):

The solution works right?

OpenStudy (the_fizicx99):

Then does are your coordinates. (4,-3)

OpenStudy (the_fizicx99):

Yes lol, did you plug it in to see if it works for the other equation? Cx It's (4,-3) I can see its 14 = 14 just by looking at it

OpenStudy (the_fizicx99):

Did you understand though? It's pointless for me to do your work if you don't understand, you'll just be lost during your exam... take it from experience <- :$ :)

OpenStudy (the_fizicx99):

@TTP did you understand :D